[Brahms-l] Shift report 20040312 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 12 2004 - 08:01:16 EST
------------------------------------------------------------
Date: Fri Mar 12 08:01:15 2004, Shift: 00:00-08:00
Supervisor: CEJ, with: FH
Log book#15 pages 86 - 86

Shift Summary:
------------------------------------------------------------
1) Beam lost at 1:15 (control failure). 
   Beam again from 5:15.

2) Pedetal runs 10924 (TPC) and 10925
   FS@10deg A861, MRS@90A1050
   10928; tr6:  67K, tr3: ??K 
   10929; tr6: 100K, tr3: 65K
   10930; tr6: 100K, tr3: 87K
   10933; still running...

3) The RunServer had to be restarted several times.
   T3-2-X1 distribution looks bad.
   No other problems reported.
   
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 12 08:01:39 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 12 2004 - 08:02:00 EST